OpenStudy (anonymous):

Solve the following equation for A: 2A/3=8+4A

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

aww thank you

OpenStudy (chris):

whoops, I meant http://openstudy.com/groups/mathematics

OpenStudy (chris):

almost sent you off into an abyss

OpenStudy (anonymous):

yea that would have been no good

OpenStudy (anonymous):

yea that bad part about it is that i cnt use a calculator its an asvab(military test) every thing i do has to on paper

OpenStudy (chris):

1. 2A/3 = 8 + 4A 2. 2A/3 x 3 = (8 + 4A) x 3 3. 2A = 24 + 12A 4. -12A = 24 5. A = 24/-12 = -2

OpenStudy (chris):

whoops. 2a/3 = 8 + 4a 2a/3 - 4a = 8 (2a - 12 a)/3 = 8 -10a = 24 a = -24/10

OpenStudy (chris):

(I did not subtract 12A from 2A correctly =))

OpenStudy (anonymous):

hahaha thank you i got the answer